# Solving General-Utility Markov Decision Processes in the Single-Trial Regime with Online Planning

### v1

@@ B.6 Proof of Theorem 3
${\boldsymbol{d}}$ $=[d(s_{0},a_{\text{include}}),d(s_{0},a_{\text{not-include}}),d(s_{1},a_{\text{include}}),d(s_{1},a_{\text{not-include}}),\ldots,$
@@
Then, we set $H\geq N$ and let $f({\boldsymbol{d}})=({\boldsymbol{n}}^{\top}{\boldsymbol{d}}-k)^{2}$, where
${\boldsymbol{n}}=\Bigg{[}\frac{1-\gamma^{H}}{1-\gamma}n_{0},0,\frac{1-\gamma^{H}}{(1-\gamma)\gamma}n_{1},0,\ldots,\frac{1-\gamma^{H}}{(1-\gamma)\gamma^{N-1}}n_{N-1},0,0,0\Bigg{]}.$
@@
For a given policy $\pi\in\Pi_{\text{NM}}^{\text{D}}$, only one trajectory $\omega\in\Omega$ has non-zero probability. The vector $\mathbf{d}^{\pi,H}_{\omega}$ associated with such a trajectory can be described as follows, for any $s_{i}\in\{0,\ldots,N-1\}$: (i) if at state $s_{i}$, $\pi$ selects $a_{\text{include}}$ then entry $\mathrm{d}^{\pi,H}_{\omega}(s_{i},a_{\text{include}})=\frac{1-\gamma}{1-\gamma^{H}}\gamma^{i}$ and $\mathrm{d}^{\pi,H}_{\omega}(s_{i},a_{\text{not-include}})=0$; (ii) if at state $s_{i}$, $\pi$ selects $a_{\text{not-include}}$ then entry $\mathrm{d}^{\pi,H}_{\omega}(s_{i},a_{\text{include}})=0$ and entry $\mathrm{d}^{\pi,H}_{\omega}(s_{i},a_{\text{not-include}})=\frac{1-\gamma}{1-\gamma^{H}}\gamma^{i}$. The action selected at $s_{N}$ is irrelevant since it does not affect the objective value. The intuition behind the GUMDP above defined is that, at each state $s_{i}$ for $i\in\{0,\ldots,N-1\}$, the policy needs to decide on whether to select action $a_{\text{include}}$ and, therefore, include term $n_{i}$ in the sum, or to select action $a_{\text{not-include}}$ and, therefore, not include term $n_{i}$ in the sum. We build the vector ${\boldsymbol{n}}$ to reflect such behavior, where each entry in ${\boldsymbol{n}}$ associated with the state $s_{i}$ and action $a_{\text{include}}$ has a normalizing constant of $\frac{1-\gamma^{H}}{(1-\gamma)\gamma^{i}}$ to account for the fact that the occupancy is discounted, as introduced in (4). Thus, it can be seen that every policy $\pi\in\Pi_{\text{NM}}^{\text{D}}$ will induce a particular trajectory $\omega\in\Omega$ with probability one and the sum of the numbers selected by the policy is given by ${\boldsymbol{n}}^{\top}{\boldsymbol{d}}^{\pi,H}_{\omega}$. Finally, the objective is such that $f({\boldsymbol{d}})=0$ if and only if the sum of the selected numbers equals $k$. The policy existence problem then asks whether there exists a policy $\pi\in\Pi_{\text{NM}}^{\text{D}}$ such that $F_{1,H}(\pi)\leq\lambda$. By setting $\lambda=0$ we are asking whether there exists a policy such that $F_{1,H}(\pi)\leq 0$. Since $f({\boldsymbol{d}})=0$ if and only if the sum of the selected numbers equals $k$, we completed our reduction from the subset problem to the policy existence problem in GUMDPs with objective $F_{1,H}$.
@@ C.1 Tasks and environments
We consider three tasks: (i) maximum state entropy exploration [Hazan et al., 2019], where the objective is to visit all state-action pairs as uniformly as possible; imitation learning [Abbeel and Ng, 2004], where the objective is to imitate a given behaviour policy; and (iii) adversarial MDPs [Rosenberg and Mansour, 2019], where an adversary player selects the cost function that yields the highest cost. We refer to Sec. 2.3 for the exact definition of the objective function for each of these tasks. We normalize all objective functions to lie in the $[0,1]$ interval. We consider two sets of environments. The first set corresponds to the illustrative GUMDPs depicted in Fig. 1, each associated with one of the tasks. The second set of environments come from the OpenAI Gym library [Brockman et al., 2016]. We consider the FrozenLake (FL), the Taxi, and the Mountaincar (MC) environments. For the MC environment, we partitioned the original state space using equally spaced bins (we consider 10 bins per dimension). For the FL, Taxi and MC environments, the task of imitation learning consists in imitating and approximately optimal policy.
@@ C.2 Experimental methodology, baselines, and hyperparameters
We perform 10 runs per experimental setting and report the 90% bootstrapped confidence interval. We let $\gamma=0.9$. We consider two baselines. The first baseline is the random policy, $\pi_{\text{Random}}$. The second baseline is the optimal policy for the infinite trials formulation (2), $\pi_{\text{Solver}}^{*}$, that we calculate by solving a constrained optimization problem with objective $f$ using the Gurobi optimizer [Gurobi Optimization, LLC, 2025]. More precisely, to compute $\pi_{\text{Solver}}^{*}$ we first solve the following optimization problem:
${\boldsymbol{d}}^{*}=\operatorname*{arg\,min}_{{\boldsymbol{d}}\in\mathcal{D}}f({\boldsymbol{d}}),$
$\mathcal{D}=\{{\boldsymbol{d}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}:d(s,a)\geq 0\;\forall s,a,\sum_{a}d(s,a)=(1-\gamma)p_{0}(s)+\gamma\sum_{s^{\prime},a}P^{a}(s|s^{\prime})d(s^{\prime},a)\;\forall s\}$

### v2

@@ B.6 Proof of Theorem 3
${\bm{d}}$ $=[d(s_{0},a_{\text{include}}),d(s_{0},a_{\text{not-include}}),d(s_{1},a_{\text{include}}),d(s_{1},a_{\text{not-include}}),\ldots,$
@@
Then, we set $H\geq N$ and let $f({\bm{d}})=({\bm{n}}^{\top}{\bm{d}}-k)^{2}$, where
${\bm{n}}=\Bigg[\frac{1-\gamma^{H}}{1-\gamma}n_{0},0,\frac{1-\gamma^{H}}{(1-\gamma)\gamma}n_{1},0,\ldots,\frac{1-\gamma^{H}}{(1-\gamma)\gamma^{N-1}}n_{N-1},0,0,0\Bigg].$
@@
For a given policy $\pi\in\Pi_{\text{NM}}^{\text{D}}$, only one trajectory $\omega\in\Omega$ has non-zero probability. The vector $\mathbf{d}^{\pi,H}_{\omega}$ associated with such a trajectory can be described as follows, for any $s_{i}\in\{0,\ldots,N-1\}$: (i) if at state $s_{i}$, $\pi$ selects $a_{\text{include}}$ then entry $\mathrm{d}^{\pi,H}_{\omega}(s_{i},a_{\text{include}})=\frac{1-\gamma}{1-\gamma^{H}}\gamma^{i}$ and $\mathrm{d}^{\pi,H}_{\omega}(s_{i},a_{\text{not-include}})=0$; (ii) if at state $s_{i}$, $\pi$ selects $a_{\text{not-include}}$ then entry $\mathrm{d}^{\pi,H}_{\omega}(s_{i},a_{\text{include}})=0$ and entry $\mathrm{d}^{\pi,H}_{\omega}(s_{i},a_{\text{not-include}})=\frac{1-\gamma}{1-\gamma^{H}}\gamma^{i}$. The action selected at $s_{N}$ is irrelevant since it does not affect the objective value. The intuition behind the GUMDP above defined is that, at each state $s_{i}$ for $i\in\{0,\ldots,N-1\}$, the policy needs to decide on whether to select action $a_{\text{include}}$ and, therefore, include term $n_{i}$ in the sum, or to select action $a_{\text{not-include}}$ and, therefore, not include term $n_{i}$ in the sum. We build the vector $\bm{n}$} to reflect such behavior, where each entry in $\bm{n}$} associated with the state $s_{i}$ and action $a_{\text{include}}$ has a normalizing constant of $\frac{1-\gamma^{H}}{(1-\gamma)\gamma^{i}}$ to account for the fact that the occupancy is discounted, as introduced in equation 4. Thus, it can be seen that every policy $\pi\in\Pi_{\text{NM}}^{\text{D}}$ will induce a particular trajectory $\omega\in\Omega$ with probability one and the sum of the numbers selected by the policy is given by ${\bm{n}}^{\top}{\bm{d}}^{\pi,H}_{\omega}$. Finally, the objective is such that $f({\bm{d}})=0$ if and only if the sum of the selected numbers equals $k$. The policy existence problem then asks whether there exists a policy $\pi\in\Pi_{\text{NM}}^{\text{D}}$ such that $F_{1,H}(\pi)\leq\lambda$. By setting $\lambda=0$ we are asking whether there exists a policy such that $F_{1,H}(\pi)\leq 0$. Since $f({\bm{d}})=0$ if and only if the sum of the selected numbers equals $k$, we completed our reduction from the subset problem to the policy existence problem in GUMDPs with objective $F_{1,H}$.
@@ C.1 Tasks and environments
We consider three tasks: (i) maximum state entropy exploration (Hazan et al., 2019), where the objective is to visit all state-action pairs as uniformly as possible; imitation learning (Abbeel and Ng, 2004), where the objective is to imitate a given behaviour policy; and (iii) adversarial MDPs (Rosenberg and Mansour, 2019), where an adversary player selects the cost function that yields the highest cost. We refer to Sec. 2.3 for the exact definition of the objective function for each of these tasks. We normalize all objective functions to lie in the $[0,1]$ interval. We consider two sets of environments. The first set corresponds to the illustrative GUMDPs depicted in Fig. 1, each associated with one of the tasks. The second set of environments come from the OpenAI Gym library (Brockman et al., 2016). We consider the FrozenLake (FL), the Taxi, and the Mountaincar (MC) environments. For the MC environment, we partitioned the original state space using equally spaced bins (we consider 10 bins per dimension). For the FL, Taxi and MC environments, the task of imitation learning consists in imitating and approximately optimal policy.
@@ C.2 Experimental methodology, baselines, and hyperparameters
We perform 10 runs per experimental setting and report the 90% bootstrapped confidence interval. We let $\gamma=0.9$. We consider two baselines. The first baseline is the random policy, $\pi_{\text{Random}}$. The second baseline is the optimal policy for the infinite trials formulation equation 2, $\pi_{\text{Solver}}^{*}$, that we calculate by solving a constrained optimization problem with objective $f$ using the Gurobi optimizer (Gurobi Optimization, LLC, 2025). More precisely, to compute $\pi_{\text{Solver}}^{*}$ we first solve the following optimization problem:
${\bm{d}}^{*}=\operatorname*{arg\,min}_{{\bm{d}}\in\mathcal{D}}f({\bm{d}}),$
$\mathcal{D}=\{{\bm{d}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}:d(s,a)\geq 0\;\forall s,a,\sum_{a}d(s,a)=(1-\gamma)p_{0}(s)+\gamma\sum_{s^{\prime},a}P^{a}(s|s^{\prime})d(s^{\prime},a)\;\forall s\}$